# 3D Reconstruction of the Human Colon from Capsule Endoscope Video

### v1

@@ 2.2 Imaging Process
Generally, we assume that the image capturing process is some mapping between 3D projective space $\mathbb{P}^{3}$ and 2D projective plane $\mathbb{P}^{2}$. Points in space, $\mathbf{X}$, are represented in homogenous coordinates $\mathbf{X}=[X,Y,Z,W]^{T}$, named world coordinates. Similarly, image points are represented as $\mathbf{x}=[x,y,w]^{T}$, named image coordinates. Here $W,w\in\mathbb{R}^{+}$ are unspecified scaling factors [7, p. 7]. For 3D points in a point cloud $\mathcal{X}$, the i’th point is denoted $\mathbf{X}_{i}$, $i=1,\cdots,N$, and its image $\mathbf{x}_{i}$. The inhomogeneous coordinates, $\tilde{\mathbf{X}}$ and $\tilde{\mathbf{x}}$, are related to the homogenous ones as [7, p. 65]
@@
With a pinhole camera the relation between a point $\mathbf{X}_{i}$ and $\mathbf{x}_{i}$ is a mapping $P:\mathbb{P}^{3}\rightarrow\mathbb{P}^{2}$. For M views of a point, $\mathbf{X}_{i}\in\mathcal{X}$, the imaging process for the j’th view is given by [7, p. 154]
@@
where $P^{j}$ is the $3\times 4$ camera matrix for the j’th view given by [7, p. 156],
@@
$R^{j}$ is a $3\times 3$ rotation matrix and $\mathbf{t}^{j}$, a $3\times 1$ translation vector. In the WCE the same camera captures all views. Therefore, the intrinsic matrix, $K$, is the same for all views [7, p. 156]
@@
where $f$ is the focal length, $p_{x},p_{y}$ is the principal point, $s$ is the skew and $m_{x},m_{y}$ is the number of pixels per unit distance in $x$- and $y$ direction. The $m_{i}$-factors in $K$ make Eq. (3) a transformation from world coordinates to pixel coordinates. For WCE’s, usually $s=0$. The other parameters can be found through a calibration procedure [7, p.226].
@@ 2.3.1 Two views and the fundamental matrix
Take Eq. (2) with $M=2$. Then any two points $\mathbf{x}_{i}^{1},\mathbf{x}_{i}^{2}$, being the images of $\mathbf{X}_{i}$ in two views, are related by the epipolar constraint [7, p.245]
@@
with $F$, the FM, a $3\times 3$ rank 2 matrix given by $F=[\mathbf{e}^{2}]_{\times}P^{2}{(P^{1})}^{+}$ [7, p.244]. Here $\mathbf{e}^{2}$ is the epipole, the image of the camera center of view 1, and ${(P^{1})}^{+}$ is the Moore-Penrose pseudoinverse of $P^{1}$. $A=[\mathbf{e}^{2}]_{\times}$ is a skew-symmetric matrix where $a_{21}=e^{2}_{3}$, $a_{31}=-e^{2}_{2}$, and $a_{32}=e^{2}_{1}$.
$F$ can be estimated from common features in two images (like SIFT features). Features are generated in the two images and matches between them are searched. With $n\geq 8$ such matches, the normalized 8-point algorithm can estimate $F$ [7, p. 282]. With significant noise in the image, outliers can be problematic, and is usually dealt with by the RANSAC [18] algorithm. With $F$ estimated, assuming that the camera center of the first view is at the world origin, the two camera matrices are given by $P^{1}=K[I|\mathbf{0}]$ and $P^{2}=[[\mathbf{e}^{2}]_{\times}F|\mathbf{e}^{2}]$ [7, p.256]. With $P^{1}$, $P^{2}$ determined, one can estimate the related 3D point, $\hat{\mathbf{X}}_{i}$, for the correspondence $\mathbf{x}_{i}^{1}\leftrightarrow\mathbf{x}_{i}^{2}$, satisfying the constraint (5), by a triangulation method [7, p. 311-313], $\hat{\mathbf{X}}_{i}=\tau(\mathbf{x}_{i}^{1},\mathbf{x}_{i}^{2},P^{1},P^{2})$.
With $\hat{\mathbf{x}}_{i}^{j}$, the projection of $\hat{\mathbf{X}}_{i}$ in the $j$-th view, one minimizes the reprojection error [7, p. 314], $C(\mathbf{x}_{i}^{1},\mathbf{x}_{i}^{2})=d(\mathbf{x}_{i}^{1},\hat{\mathbf{x}}_{i}^{1})^{2}+d(\mathbf{x}_{i}^{2},\hat{\mathbf{x}}_{i}^{2})^{2}$, subject to (5), with $d(\cdot,\cdot)$ some distance measure.
With $K$ known, one can estimate the Essential Matrix (EM) $E=K^{T}FK$ [7, p. 257], instead of the FM, which is simpler to compute. With $K$ and $F$ (or $E$) known, the 3D scene can be reconstructed up to a similarity transform [7, p. 272-273], i.e, a Euclidean reconstruction up to an unknown scaling factor. The exception is the degenerate case, which can occur when the camera centers and $\mathbf{X}_{i}$ are co-linear (practically close to co-linear). Under pure rotation about the camera center, the degenerate case $F=0$ occurs. For WCE, degeneracies may occur due to both of these cases. To obtain a metric reconstruction, information about the size of some object in the scene is needed. An effort dealing with this for WCE application is [19].
@@ 2.3.2 Multiple Views and Bundle Adjustment (BA)
For $M>4$ views, the problem has to be dealt with numerically through bundle adjustment (BA), which is a minimization problem on the form [7, p. 434],
@@
with $d(\cdot,\cdot)$, some distance measure, typically Euclidean norm. That is, BA is the reprojection error over all views and 3D points. BA requires a good initial estimate, $\hat{P}^{j},\hat{\mathbf{X}}_{i}$, of camera poses and 3D points, typically obtained by computing the FM sequentially over pairs of neighboring images for all views in the sequence [7, p.453]. BA is costly to compute for large $M$ [7, p. 435], a problem solved by ORB-SLAM.
@@ 2.5 Poisson Surface Reconstruction
The implementation is built around an octree, a tree where each node, except leaf nodes, has 8 children, and all samples, $\tilde{\mathbf{X}}_{i}\in\tilde{\mathcal{X}}$, are located in leaf nodes. The space containing the object $M$ is subdivided into 8 cubes at each level of the tree so that finer resolution is obtained at deeper levels. To each leaf node, $o$, there is an associated function $F_{o}$, typically approximating a Gaussian filter, whose width is adapted to the local sample resolution [15]. With $D_{t}$, the depth of the octree, the relevant space will be divided into $8^{D_{t}}$ cubes. This leads to a multiresolution structure where points in sparse neighborhoods are located in leaf nodes high up in the tree, whereas points in dense neighborhoods are located further down. Details on the implementation can be found in [15, Sections 4.3-4.5].
@@ 2.6 Point Cloud Alignment and Error Computation
where $\mathring{q}^{\ast}$ is the conjugate of $\mathring{q}$ and $\mathring{r}=\mathbf{i}r_{x}+\mathbf{j}r_{y}+\mathbf{k}r_{z}$ is the imaginary quaternion. How to determine the max of (13) and the corresponding algorithm is given in [22, pp. 635-636].
@@ 3.1 Structure from motion
SURF features are used to estimate $E$ (and thereby $P^{i}$) and the 3D points, whereas eigenfeatures are used to compute dense point clouds once $E$ is known. For $M=6$ views, an initial reconstruction is made by sequentially computing $E$ for pairs of consecutive frames (as in Algorithm 18.3 in [7, p. 453]) followed by BA. All relevant computation and estimation methods for our purposes are found in Matlab’s computer vision toolbox [^3] https://se.mathworks.com/products/computer-vision.html (10/11-21).

### v2

@@ 1 Introduction
NOTE: This is a preliminary version of this paper. The final complete paper is to be found in IEEE ACCESS, Volume 13, DOI: 10.1109/ACCESS.2025.3587596 [1].
@@ 2.2 Imaging Process
Generally, we assume that the image capturing process is some mapping between 3D projective space $\mathbb{P}^{3}$ and 2D projective plane $\mathbb{P}^{2}$. Points in space, $\mathbf{X}$, are represented in homogenous coordinates $\mathbf{X}=[X,Y,Z,W]^{T}$, named world coordinates. Similarly, image points are represented as $\mathbf{x}=[x,y,w]^{T}$, named image coordinates. Here $W,w\in\mathbb{R}^{+}$ are unspecified scaling factors [8, p. 7]. For 3D points in a point cloud $\mathcal{X}$, the i’th point is denoted $\mathbf{X}_{i}$, $i=1,\cdots,N$, and its image $\mathbf{x}_{i}$. The inhomogeneous coordinates, $\tilde{\mathbf{X}}$ and $\tilde{\mathbf{x}}$, are related to the homogenous ones as [8, p. 65]
@@
With a pinhole camera the relation between a point $\mathbf{X}_{i}$ and $\mathbf{x}_{i}$ is a mapping $P:\mathbb{P}^{3}\rightarrow\mathbb{P}^{2}$. For M views of a point, $\mathbf{X}_{i}\in\mathcal{X}$, the imaging process for the j’th view is given by [8, p. 154]
@@
where $P^{j}$ is the $3\times 4$ camera matrix for the j’th view given by [8, p. 156],
@@
$R^{j}$ is a $3\times 3$ rotation matrix and $\mathbf{t}^{j}$, a $3\times 1$ translation vector. In the WCE the same camera captures all views. Therefore, the intrinsic matrix, $K$, is the same for all views [8, p. 156]
@@
where $f$ is the focal length, $p_{x},p_{y}$ is the principal point, $s$ is the skew and $m_{x},m_{y}$ is the number of pixels per unit distance in $x$- and $y$ direction. The $m_{i}$-factors in $K$ make Eq. (3) a transformation from world coordinates to pixel coordinates. For WCE’s, usually $s=0$. The other parameters can be found through a calibration procedure [8, p.226].
@@ 2.3.1 Two views and the fundamental matrix
Take Eq. (2) with $M=2$. Then any two points $\mathbf{x}_{i}^{1},\mathbf{x}_{i}^{2}$, being the images of $\mathbf{X}_{i}$ in two views, are related by the epipolar constraint [8, p.245]
@@
with $F$, the FM, a $3\times 3$ rank 2 matrix given by $F=[\mathbf{e}^{2}]_{\times}P^{2}{(P^{1})}^{+}$ [8, p.244]. Here $\mathbf{e}^{2}$ is the epipole, the image of the camera center of view 1, and ${(P^{1})}^{+}$ is the Moore-Penrose pseudoinverse of $P^{1}$. $A=[\mathbf{e}^{2}]_{\times}$ is a skew-symmetric matrix where $a_{21}=e^{2}_{3}$, $a_{31}=-e^{2}_{2}$, and $a_{32}=e^{2}_{1}$.
$F$ can be estimated from common features in two images (like SIFT features). Features are generated in the two images and matches between them are searched. With $n\geq 8$ such matches, the normalized 8-point algorithm can estimate $F$ [8, p. 282]. With significant noise in the image, outliers can be problematic, and is usually dealt with by the RANSAC [19] algorithm. With $F$ estimated, assuming that the camera center of the first view is at the world origin, the two camera matrices are given by $P^{1}=K[I|\mathbf{0}]$ and $P^{2}=[[\mathbf{e}^{2}]_{\times}F|\mathbf{e}^{2}]$ [8, p.256]. With $P^{1}$, $P^{2}$ determined, one can estimate the related 3D point, $\hat{\mathbf{X}}_{i}$, for the correspondence $\mathbf{x}_{i}^{1}\leftrightarrow\mathbf{x}_{i}^{2}$, satisfying the constraint (5), by a triangulation method [8, p. 311-313], $\hat{\mathbf{X}}_{i}=\tau(\mathbf{x}_{i}^{1},\mathbf{x}_{i}^{2},P^{1},P^{2})$.
With $\hat{\mathbf{x}}_{i}^{j}$, the projection of $\hat{\mathbf{X}}_{i}$ in the $j$-th view, one minimizes the reprojection error [8, p. 314], $C(\mathbf{x}_{i}^{1},\mathbf{x}_{i}^{2})=d(\mathbf{x}_{i}^{1},\hat{\mathbf{x}}_{i}^{1})^{2}+d(\mathbf{x}_{i}^{2},\hat{\mathbf{x}}_{i}^{2})^{2}$, subject to (5), with $d(\cdot,\cdot)$ some distance measure.
With $K$ known, one can estimate the Essential Matrix (EM) $E=K^{T}FK$ [8, p. 257], instead of the FM, which is simpler to compute. With $K$ and $F$ (or $E$) known, the 3D scene can be reconstructed up to a similarity transform [8, p. 272-273], i.e, a Euclidean reconstruction up to an unknown scaling factor. The exception is the degenerate case, which can occur when the camera centers and $\mathbf{X}_{i}$ are co-linear (practically close to co-linear). Under pure rotation about the camera center, the degenerate case $F=0$ occurs. For WCE, degeneracies may occur due to both of these cases. To obtain a metric reconstruction, information about the size of some object in the scene is needed. An effort dealing with this for WCE application is [20].
@@ 2.3.2 Multiple Views and Bundle Adjustment (BA)
For $M>4$ views, the problem has to be dealt with numerically through bundle adjustment (BA), which is a minimization problem on the form [8, p. 434],
@@
with $d(\cdot,\cdot)$, some distance measure, typically Euclidean norm. That is, BA is the reprojection error over all views and 3D points. BA requires a good initial estimate, $\hat{P}^{j},\hat{\mathbf{X}}_{i}$, of camera poses and 3D points, typically obtained by computing the FM sequentially over pairs of neighboring images for all views in the sequence [8, p.453]. BA is costly to compute for large $M$ [8, p. 435], a problem solved by ORB-SLAM.
@@ 2.5 Poisson Surface Reconstruction
The implementation is built around an octree, a tree where each node, except leaf nodes, has 8 children, and all samples, $\tilde{\mathbf{X}}_{i}\in\tilde{\mathcal{X}}$, are located in leaf nodes. The space containing the object $M$ is subdivided into 8 cubes at each level of the tree so that finer resolution is obtained at deeper levels. To each leaf node, $o$, there is an associated function $F_{o}$, typically approximating a Gaussian filter, whose width is adapted to the local sample resolution [16]. With $D_{t}$, the depth of the octree, the relevant space will be divided into $8^{D_{t}}$ cubes. This leads to a multiresolution structure where points in sparse neighborhoods are located in leaf nodes high up in the tree, whereas points in dense neighborhoods are located further down. Details on the implementation can be found in [16, Sections 4.3-4.5].
@@ 2.6 Point Cloud Alignment and Error Computation
where $\mathring{q}^{\ast}$ is the conjugate of $\mathring{q}$ and $\mathring{r}=\mathbf{i}r_{x}+\mathbf{j}r_{y}+\mathbf{k}r_{z}$ is the imaginary quaternion. How to determine the max of (13) and the corresponding algorithm is given in [23, pp. 635-636].
@@ 3.1 Structure from motion
SURF features are used to estimate $E$ (and thereby $P^{i}$) and the 3D points, whereas eigenfeatures are used to compute dense point clouds once $E$ is known. For $M=6$ views, an initial reconstruction is made by sequentially computing $E$ for pairs of consecutive frames (as in Algorithm 18.3 in [8, p. 453]) followed by BA. All relevant computation and estimation methods for our purposes are found in Matlab’s computer vision toolbox [^3] https://se.mathworks.com/products/computer-vision.html (10/11-21).